OpenStudy (mortonsalt):

Hello! I'm a little confused on how to graph an exponential function like y = 3x^-1. Does anyone know any videos that could help me out?

OpenStudy (unklerhaukus):

that's not an exponential function, that is a hyperbola

OpenStudy (mortonsalt):

Well, there's my problem right there. Hahaha! Thank you.

OpenStudy (anonymous):

If you want to know the shape of that function it's just

OpenStudy (anonymous):

Kind of

OpenStudy (mortonsalt):

Thanks! It's a lot easier for me to find now.

OpenStudy (unklerhaukus):

exponential functions are like \[y = b^x\] hyperbolas are like \[y = \frac ax\]

OpenStudy (unklerhaukus):

The vertical asymptote of the hyperbola is when the denominator term approaches zero

OpenStudy (mortonsalt):

I kind of looked at them as y = 3x^-1. Which is obviously very wrong. hahaha Thank you very much